A jet lands on an aircraft carrier at 63 m/s.

a) What is its acceleration (assumed constant) if it stops in 2.0 s due to an

arresting cable that snags the airplane and brings it to a stop?

b) If the plane touches down at position xo=0m, what is the final position of

the plane?


1
Expert's answer
2021-04-08T09:02:00-0400

(a)


"v=v_0-at\to a=v_0/t=63/2=31.5\ (m/s^2)" . Answer


(b)


"x=v_0^2/(2a)=63^2/(2\cdot31.5)=63\ (m)" . Answer